OpenStudy (kanard):

A square rug covers 79 square feet of floor. What is the approximate length of one side of the rug? (Approximate to the nearest hundredth foot.) 8.86 feet 8.87 feet 8.88 feet 8.89 feet

OpenStudy (blow_pop2000):

──────▄▀▀▄────────────────▄▀▀▄──── ─────▐▒▒▒▒▌──────────────▌▒▒▒▒▌─── ─────▌▒▒▒▒▐─────────────▐▒▒▒▒▒▐─── ────▐▒▒▒▒▒▒▌─▄▄▄▀▀▀▀▄▄▄─▌▒▒▒▒▒▒▌── ───▄▌▒▒▒▒▒▒▒▀▒▒▒▒▒▒▒▒▒▒▀▒▒▒▒▒▒▐─── ─▄▀▒▐▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▌─── ▐▒▒▒▌▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▐─── ▌▒▒▌▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▌── ▒▒▐▒▒▒▒▒▒▒▒▒▄▀▀▀▀▄▒▒▒▒▒▄▀▀▀▀▄▒▒▐── ▒▒▌▒▒▒▒▒▒▒▒▐▌─▄▄─▐▌▒▒▒▐▌─▄▄─▐▌▒▒▌─ ▒▐▒▒▒▒▒▒▒▒▒▐▌▐█▄▌▐▌▒▒▒▐▌▐█▄▌▐▌▒▒▐─ ▒▌▒▒▒▒▒▒▒▒▒▐▌─▀▀─▐▌▒▒▒▐▌─▀▀─▐▌▒▒▒▌ ▒▌▒▒▒▒▒▒▒▒▒▒▀▄▄▄▄▀▒▒▒▒▒▀▄▄▄▄▀▒▒▒▒▐ ▒▌▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▄▄▄▒▒▒▒▒▒▒▒▒▒▒▐ ▒▌▒▒▒▒▒▒▒▒▒▒▒▒▀▒▀▒▒▒▀▒▒▒▀▒▀▒▒▒▒▒▒▐ ▒▌▒▒▒▒▒▒▒▒▒▒▒▒▒▀▒▒▒▄▀▄▒▒▒▀▒▒▒▒▒▒▒▐ ▒▐▒▒▒▒▒▒▒▒▒▒▀▄▒▒▒▄▀▒▒▒▀▄▒▒▒▄▀▒▒▒▒▐ ▒▓▌▒▒▒▒▒▒▒▒▒▒▒▀▀▀▒▒▒▒▒▒▒▀▀▀▒▒▒▒▒▒▐ ▒▓▓▌▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▌ ▒▒▓▐▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▌─ ▒▒▓▓▀▀▄▄▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▐── ▒▒▒▓▓▓▓▓▀▀▄▄▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▄▄▀▀▒▌─ ▒▒▒▒▒▓▓▓▓▓▓▓▀▀▀▀▀▀▀▀▀▀▀▀▀▀▀▒▒▒▒▒▐─ ▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▌ ▒▒▒▒▒▒▒█▒█▒█▀▒█▀█▒█▒▒▒█▒█▒█▒▒▒▒▒▒▐ ▒▒▒▒▒▒▒█▀█▒█▀▒█▄█▒▀█▒█▀▒▀▀█▒▒▒▒▒▒▐ ▒▒▒▒▒▒▒▀▒▀▒▀▀▒▀▒▀▒▒▒▀▒▒▒▀▀▀▒▒▒▒▒▒▐ █▀▄▒█▀▄▒█▀▒█▀█▒▀█▀▒█▒█▒█▒█▄▒█▒▄▀▀▐ █▀▄▒█▀▄▒█▀▒█▄█▒▒█▒▒█▀█▒█▒█▀██▒█▒█▐ ▀▀▒▒▀▒▀▒▀▀▒▀▒▀▒▒▀▒▒▀▒▀▒▀▒▀▒▒▀▒▒▀▀▐ ▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▒▐

OpenStudy (kanard):

^ wtf..

OpenStudy (kanard):

@AlexandervonHumboldt2 Last one of the night could you help me ;-;

OpenStudy (kanard):

@dan815 @ganeshie8 @pooja195

OpenStudy (anonymous):

what is the equation for the area of a square?

OpenStudy (anonymous):

@kanard

OpenStudy (anonymous):

i only have a few minutes so we gotta do this quick haha

OpenStudy (anonymous):

A=s^2 so if the area is given plug it in and solve for s: 79=s^2 \[\sqrt{79}=\sqrt{s^2}\] so \[s \approx 8.88\]